A tank can be filled by two pipes separately in 10 and 15 minutes respectively. When a third pipe is used simultaneously with the first two pipes, the tank can be filled in 4 minutes. How long would it take for the third pipe alone to fill the tank?
Answer by stanbon(75887)

A tank can be filled by two pipes separately in 10 and 15 minutes respectively. When a third pipe is used simultaneously with the first two pipes, the tank can be filled in 4 minutes. How long would it take for the third pipe alone to fill the tank?
--------
1st pipe rate: 1/10 job/minute
2nd pipe rate: 1/15 job/minute
3rd pipe rate: 1/x job/min
Together rate: 1/4 job/min
-----
Equation:
rate + rate + rate = together rate::
1/10 + 1/15 + 1/x = 1/4
------------
6x + 4x + 60 = 15x
5x = 60
x = 12 minutes (time for the 3rd pipe to fill the tank)
============
Cheers,
Stan H.
=============